via this auction

Service work completed:
Tightened loose key contact actuators
Fixed stuck key contact (low C)
Adjusted key heights
Replaced damaged/bad slide pots (HPF cutoff frequency 1M/A, both frequency coarse 100kB, & VCO-1 level 100kA)
Calibrated supply voltages, VCOs, transpose, & VCF
Performed burn-in test
Tightened loose hardware"
No comments:
Post a Comment
Note: comments that insult people will be removed. Critique on gear is allowed. Do not ask if listings are still available. Click through auction links to check yourself. Posts and pics remain for historical purposes. To reduce spam, comments for posts older than one week are not displayed until approved (usually same day).